# Automated radiotherapy treatment planning guided by GPT-4Vision

### v2

@@ 4 Discussion
The proposed GPT-RadPlan represents an advancement in the field of radiation oncology, offering a fully automated treatment planning framework that leverages the capabilities of state-of-the-art multi-modal LLMs, such as OpenAI GPT-4V. Its ability to operate without the need for domain-specific datasets or extensive training facilitates integrating GPT-RadPlan into various clinical workflows, without requiring any adjustment to current clinical software and optimization algorithms. This increased efficiency means that clinicians can deliver personalized treatment plans much faster, reducing patient wait times and allowing healthcare facilities and increasing clinical throughput. Moreover, being clinical software agnostic, GPT-RadPlan can be readily adopted across diverse treatment protocols and institutions, underscoring its broad applicability. To our knowledge, this is the first time that an automated treatment planning system based on LLMs has been proposed and demonstrated to be feasible for direct application to clinical environments without requiring any extensive model training.
@@
GPT-RadPlan has two advantages over any other learning-based method. First, it does not require a large dataset or time-consuming training, leveraging LLM’s in-context learning abilities instead with only a handful of approved clinical plans. Moreover, users can specify clinical preferences and protocols with human language, making GPT-RadPlan a versatile tool that can generalize across disease sites, protocol requirements and metrics. As such, GPT-RadPlan can optimize plans to directly satisfy prescribed clinical criteria as defined by radiation oncologists, therefore avoiding the use of proxy metrics or crafted wish-lists. The reduced need for human interventions during planning further enhances clinical workflow and improves the allocation of human resources,freeing radiation oncologists and dosimetrists to focus on other critical aspects of patient care.
@@
GPT-RadPlan also offers several advantages in comparison to previous RL auto-planning methods Shen et al. (2020). In general, deep RL algorithms approximate policy functions and take actions in a ’black-box’ manner, where decisions are hardly interpretable. In contrast, GPT-RadPlan reflects on its previous results, suggesting improvements and acting based on these reflections. As a result, GPT-RadPlan actions (i.e., weight adjustments) always have an explanation. Moreover, compared to the previous RL studies requiring an entire week to train an RL agent for a specific disease site and protocol, GPT-RadPlan massively reduces computation times, being directly applicable after only minor prompt modifications. Lastly, unlike RL agents requiring a reward that hardly captures all clinical objectives and human preferences, the proposed GPT-RadPlan agents directly evaluate plans based on the entire prescribed protocol fully capturing the physician’s intent.
@@
As demonstrated in the results, the plans generated by GPT-RadPlan uniformly cover PTV targets and spare OARs, providing a high level of consistency and quality that should contribute to better patient outcomes. After multiple iterations, GPT-RadPlan can provide plans that satisfy all the criteria in the clinical protocol for prostate cases, matching the PTV coverage requirements while always irradiating OARs with lower mean doses and DVH constraints than those indicated in the protocol. For head & neck plans, GPT-RadPlan’s plans cannot satisfy all the clinical requirements simultaneously, as is often observed in clinical practice. This involves conflicting objectives, such as sparing two different organs or prioritizing the coverage of the PTVs rather than sparing a single OAR (e.g., the larynx). Since the planner prioritizes dose objectives based on the order specified in the prompt, GPT-RadPlan could, in principle, allow for the exploration of the Pareto frontier by easily modifying these prompt priorities.
As possible next steps, instead of providing a few manually selected patients as a reference, GPT-RadPlan could be coupled to retrieval-augmented generation (RAG) that retrieves the most similar plans in a database to be used for assistance Huang et al. (2023); S. Liu, Kaku, et al. (2022). Future work could also explore GPT-RadPlan applications to proton treatments in the presence of setup, range or anatomical uncertainties Pastor-Serrano et al. (2021, 2023), expanding the model capabilities to adjust robustness settings in addition to objective weights. Additionally, since one limitation of GPT-RadPlan is its lack of direct alignment with human preferences, subsequent studies could further improve model performance by manually collecting radiation oncologists’ plan evaluations and using these to fine-tune the LLMs and calibrating the LLMs Y. Liu et al. (2024); S. Liu, Zhang, et al. (2022); S. Liu, Kaku, et al. (2022), or using RT fine-tuned models to assist in the evaluation task Z. Liu et al. (2023).
@@ 5 Conclusion
6 Acknowledgement
This research was supported in part by the National Institutes of Health (NIH) grant 1R01CA256890 and 1R01CA275772.
7 Conflict of interest
O. Pastor-Serrano, S. Liu, J. Zou and L. Xing are co-inventors on a patent application based on this work.

### v3

@@ 4 Discussion
The proposed GPT-RadPlan represents an advancement in the field of radiation oncology, offering a fully automated treatment planning framework that leverages the capabilities of state-of-the-art multi-modal LLMs, such as OpenAI GPT-4V. Its ability to operate without the need for domain-specific datasets or extensive training potentially facilitates integrating GPT-RadPlan into various clinical workflows without requiring any adjustment to current clinical software and optimization algorithms. Notably, the code we use is identical for both prostate and head & neck patients, with the only variation being in the input configuration file, which specifies the location of reference cases, the physician’s intent in text form, and a few other machine parameters. Moreover, being clinical software agnostic, GPT-RadPlan has the potential to be readily adopted across diverse disease sites and institutions, underscoring its broad applicability. To our knowledge, this is the first time that an automated treatment planning system based on LLMs has been proposed and demonstrated to be feasible for direct application to clinical environments without requiring any extensive model training.
@@
GPT-RadPlan has two advantages over any other learning-based method. First, it does not require a large dataset or time-consuming training, leveraging LLM’s in-context learning abilities instead with only a handful of approved clinical plans. Second, it uses all the available information to evaluate and improve plans, including dose distribution, DVH, clinical preferences (as text) and quantitative metrics. Moreover, users can potentially specify clinical preferences for a specific patient using human language (though we did not do it in our experiments, making GPT-RadPlan a versatile tool that can potentially be adapted to multiple disease sites and requirements using the same code and exclusively changing the prompt. As such, GPT-RadPlan can optimize plans to directly satisfy prescribed clinical criteria as defined by radiation oncologists in the physician’s intent, therefore avoiding the use of proxy metrics or crafted wish-lists. The reduced need for human interventions during planning further enhances clinical workflow and improves the allocation of human resources, freeing radiation oncologists and dosimetrists to focus on other critical aspects of patient care.
@@
GPT-RadPlan also offers several advantages in comparison to previous RL auto-planning methods [24]. In general, deep RL algorithms approximate policy functions and take actions in a ’black-box’ manner, where decisions are hardly interpretable. In contrast, GPT-RadPlan reflects on its previous results, suggesting improvements and acting based on these reflections. As a result, GPT-RadPlan actions (i.e., weight adjustments) always have an explanation. Moreover, compared to the previous RL studies requiring an entire week to train an RL agent for a specific disease site and protocol, GPT-RadPlan massively reduces computation times, being directly applicable after only minor prompt modifications. Lastly, unlike RL agents requiring a reward that hardly captures all clinical objectives and human preferences, the proposed GPT-RadPlan agents directly evaluate plans based on the entire prescribed protocol fully capturing the physician’s intent.
@@
As demonstrated in the results, the plans generated by GPT-RadPlan uniformly cover PTV targets and spare OARs, providing a high level of consistency and quality that should contribute to better patient outcomes. After multiple iterations, GPT-RadPlan can provide plans that satisfy all the criteria in the clinical protocol for prostate cases, matching the PTV coverage requirements while always irradiating OARs with lower mean doses and DVH constraints than those indicated in the protocol. For head & neck plans, GPT-RadPlan’s plans cannot satisfy all the clinical requirements simultaneously, as is also often observed in clinical practice. This involves conflicting objectives, such as sparing two different organs or prioritizing the coverage of the PTVs rather than sparing a single OAR (e.g., the larynx). Since the AI planner prioritizes dose objectives based on the order specified in the prompt, GPT-RadPlan could, in principle, allow for the exploration of the Pareto frontier by easily modifying these prompt priorities.
Future steps
As possible next steps, instead of providing a few manually selected patients as a reference, GPT-RadPlan could be coupled to retrieval-augmented generation (RAG) that retrieves the most similar plans in a database to be used for assistance [10, 37]. Future work could also explore GPT-RadPlan applications to proton treatments in the presence of setup, range or anatomical uncertainties [38, 39], expanding the model capabilities as a treament planninf copilot to automatically select robustness settings or the anatomies or scenarios included during robust treatment planning. This work proposed novel method to automate treatment planning using pre-trained LLMs that shows promising results in just two disease sites, introducing a new line of research that explores the advantages of applying LLMs to assist human clinicians. To demonstrate generalization across disease sites, future work should focus on the clinical evaluation of GPT-RadPlan in more diverse clinical settings and benchmarks.
Limitations
This paper presents the novel idea of using LLMs as a copilot during treatment planning to mimic human planners decisions, but the TPS used to obtain clinical plans (Eclipse) is different from the one that GPT-RadPlan is coupled to. Although the objectives and problem formulation are the same, subsequent studies must compare GPT-RadPlan with human clinical planners using the same TPS.
One limitation of GPT-RadPlan is its lack of direct alignment with physicians’ preferences through model fine-tuning as we directly use the GPT-4V, which is not specifically trained for the task. While it is true that OpenAI trained their models on large-scale datasets that may include detailed clinical information, we do not supply any specific information about optimal planning choices for the patients. The model can only be made aware of clinical preferences via the physician’s intent in the prompt, which limits the ability to specify trade-off preferences that could significantly differ between areas, e.g., the larynx and oropharynx. Subsequent studies could further improve model performance by manually collecting radiation oncologists’ plan evaluations and using these to fine-tune the LLMs and calibrating the LLMs [40, 41, 37], or using RT fine-tuned models to assist in the evaluation task [42]. Due to the dependence on input clinical protocols and reference cases, ambiguous or inadequate input could result in GPT-RadPlan generating suboptimal or potentially unsafe treatment plans. Furthermore, the use of non-representative reference cases may introduce bias, limiting the generalization of the model across diverse patient populations or institutions. While GPT-RadPlan provides interpretable feedback for parameter adjustments, its underlying decision-making remains partially opaque, which could hinder clinicians’ ability to fully understand or trust its recommendations.
To mitigate these risks, continuous physician supervision is crucial to ensure adherence to safety and quality standards. Regular calibration and testing of GPT-RadPlan using real-world feedback can help maintain its reliability and adapt it to diverse clinical settings. In addition, incorporating automated validation mechanisms to flag outlier plans or deviations from clinical guidelines can serve as a robust safety net.
Future efforts should focus on fine-tuning GPT-RadPlan with localized datasets to enhance its adaptability, developing explainability frameworks for greater transparency, and integrating uncertainty quantification methods to identify high-risk predictions. These steps will further improve the reliability and clinical utility of GPT-RadPlan.
Finally, a limitation of GPT-RadPlan is its inability to match the Pareto optimality guarantees of multicriteria optimization methods [13, 12], which generate a Pareto optimal set of plans where no single objective (like tumor coverage or organ sparing) can be improved without compromising another. Trial-and-error methods such GPT-RadPlan do not explore the Pareto front but rather converge to a solution that is clinically acceptable based on physician intent and specified preferences. While, as demonstrated in the results, this final plan can perform well, GPT-RadPlan does not provide Pareto optimality guarantees or a clear view of the planning trade-offs.
@@ 5 Conclusion
Appendix A Prompts
Radiotherapy treatment planning involves an evaluator that evaluates the inverse optimized plan, and a planner that produces the parameters for the next iteration. We adopt multimodal LLMs to evaluate and produce new parameters by prompting them to perform the corresponding tasks. Specifically, for the evaluation, a multimodal LLM is used to compare each protocol with the current plan and produce the final assessment. Similarly, we prompt a multimodal LLM to produce new parameters based on the plan in the current iteration as well as feedback from the evaluator.
Evaluator prompt Please act as an impartial and objective professional radiation oncologist. Your job is to evaluate the quality of radiation therapy treatment plans based on their dose volume histogram. You should consider the following protocol: <Clinical Protocol> Here is the dose-volume histograms of the candidate plans for evaluation; each entry in the dose-volume histograms (DVH) table indicates the percentage of volume receiving a dose higher than a certain Gy (specified in the first column). <DVH table> We also provide the statistics of the above DVH table for ease of evaluation. <DVH statistics> You should also evaluate the plan based on the dose distribution, considering the anatomy of the patient. For your evaluation, you should consider the following factors: Homogeneity and conformity: A good plan should have a conformal and homogeneous dose (red in the image) within the PTV. There should be no hot spots or cold spots. Dose spillage. A good plan should NOT have high (red) dose spillage outside of the PTV. The does distribution must show a sharp gradient outside the PTV/target. <Images> Now, based on the protocol, the images and the DVH, please evaluate the plans. Avoid any positional biases and ensure that the order in which the responses are presented does not influence your decision. Be as objective as possible. Your answer must: 1. (Evidence) Extract the corresponding entries from the DVH table based on the factors to consider. 2. (Interpretation) Based on the extracted entries from DVH table, interpret them as: {percentage}of volume encompassing the dose more than {dose} Gy for {organ}. 3. (Analysis) Combine the interpretation and the factors, analyze whether each of the numerical factors are met. Based on the analysis on DVH tables, you need to produce a final evaluation. When all factors are satisfied, you should always answer no improvement is needed. If improvements are required, please provide suggestions on where to improve. Please ensure to follow the format below to return the evaluation results: <FINAL> Decision: [The plan does/doesn’t need to be improved.] Reasons: [list all factors that are not satisfied with detailed reasons, e.g. protocol X on (PTV/rectum/bladder/fh/body) is not satisfied because …] </FINAL> The final answer in the end must strictly follow the format above.
Planner prompt Please act as a professional radiation oncologist who is impartial and objective. Your job is to provide a set of parameters to optimize the treatment plan. Treatment planning involves a balancing act between competing clinical goals. The desirable dose distribution is determined by an optimization algorithm minimizing a cost function with multiple competing objectives. Each objective is the difference between the current dose and the prescribed dose of a certain organ (PTV, brainstem, etc.). The parameters decide the importance of each objective. You are provided with a set of good reference plans, the DVH, and their corresponding parameters, to learn what the parameters of good plans should look like. Reference treatment plan DVH: <Reference plan DVH> Reference treatment plan parameters: <Reference plan parameters> In the following, you are provided with the dose volume histograms (DVHs) of the treatment plans in the previous iterations as well as the parameter weights used to generate the plans. Here are suggestions from the evaluator on improving the historical treatment plan: <suggestions on improving the plan> Now, based on the historical treatment plan as well as the suggestion on what to improve, you need to generate a new set of parameters to rebalance the objectives of the optimization: Please ensure to conclude with the following one-line dictionary format (without name): "PTV70": <weight>, "PTV56": <weight>, "PTV52": <weight>, "Brainstem": <weight>, "SpinalCord": <weight>, "ParotidLeft": <weight>, "ParotidRight": <weight>, "OralCavity": <weight>, "Larynx": <weight>, "Body": <weight> The final answer in the end must strictly follow this format in one single line to get a valid Python dictionary that I can directly parse it using Python.
Appendix B Additional results
In this section, we present metrics and figures to further evaluate the OAR sparing capabilities of GPT-RadPlan with respect to clinical plans (Section B.1). We also include additional plots comparing the proposed method to the non-clinical Autoplan baseline in Section B.2.
B.1 Clinical plan comparison
As observed in Table 1, GPT-RadPlan delivers a lower mean dose to the PTV that is closer to prescription, especially in prostate treatments. To discard lower total average doses as a potential cause of observing lower OAR mean doses and metrics, we have rescaled the GPT-RadPlan plan doses so that the mean PTV dose matches that of clinical plans. Based on these rescaled doses, Table 3 shows the recalculated OAR metrics. These are the same metrics and calculations as in Table 2. With an average 4 Gy difference for prostate plans and 6 Gy difference for head & neck plans, these results confirm GPT-RadPlan’s OAR sparing capabilities.
Table 3: OAR dose metrics after dose scaling. Multiple dose metrics capturing OAR sparing in prostate and head & neck plans are presented, after scaling the GPT-RadPlan doses so that the mean PTV doses match those of the clinical plans. These include the mean dose, the $D_{5}$ and $D_{50}$ (maximum common dose that 5% or 50% of the volume receives), and the $V_{15}$ and $V_{30}$ (representing which percentage of the volume receives 15 and 30 Gy, respectively). For all the metrics, we include the average values across plans and the standard deviation in brackets. In all cases, lower values are preferred.
Prostate plans
Bladder Clinical 24.48 (9.30) 62.14 (10.79) 18.87 (12.12) 60.85 (29.52) 34.45 (15.28)
GPT-RadPlan 17.41 (8.34) 61.42 (14.77) 9.41 (8.85) 36.41 (17.52) 23.13 (13.46)
Rectum Clinical 26.45 (5.64) 63.64 (8.82) 22.20 (5.71) 71.91 (14.92) 32.87 (10.54)
GPT-RadPlan 22.27 (6.68) 60.18 (14.13) 17.44 (8.06) 51.52 (20.39) 28.21 (16.64)
Head & neck plans
Brainstem Clinical 12.17 (4.29) 30.95 (4.78) 8.34 (5.75) 33.36 (14.10) 10.03 (11.03)
GPT-RadPlan 10.8 (4.82) 26.57 (4.65) 7.98 (7.41) 32.54 (15.84) 0.03 (0.1)
Larynx Clinical 42.80 (11.03) 69.95 (4.30) 39.85 (14.87) 97.44 (3.24) 64.06 (18.18)
GPT-RadPlan 33.78 (14.36) 69.82 (5.77) 26.71 (21.6) 63.42 (26.83) 42.04 (24.98)
Oral Cavity Clinical 34.86 (4.15) 68.07 (8.83) 33.62 (12.29) 92.96 (5.28) 50.00 (11.01)
GPT-RadPlan 25.12 (8.45) 65.58 (14.99) 18.04 (8.26) 53.74 (25.5) 27.59 (13.65)
Parotid Left Clinical 35.10 (13.93) 63.18 (10.85) 33.50 (21.21) 71.64 (20.97) 52.80 (24.12)
GPT-RadPlan 23.99 (12.54) 57.08 (15.96) 19.51 (18.42) 42.5 (20.52) 28.18 (20.87)
Parotid Right Clinical 34.32 (10.67) 61.69 (11.30) 32.22 (13.83) 79.00 (21.05) 52.65 (20.65)
GPT-RadPlan 21.41 (7.74) 57.8 (15.47) 13.51 (7.41) 41.36 (19.55) 25.33 (13.72)
Spinal Cord Clinical 17.36 (6.32) 33.88 (5.15) 18.84 (13.24) 55.32 (13.69) 24.83 (25.69)
GPT-RadPlan 16.07 (3.93) 29.77 (0.3) 17.56 (12.37) 55.12 (14.1) 1.33 (3.22)
Furthermore, to help visualize the individual plan-by-plan differences between GPT-RadPlan and clinical plans from Figure 4, we have included additional Figure 6 and Figure 7, where each bar shows the same relative error for every OAR and plan in the prostate and head & neck cohorts, respectively. Positive values indicate higher clinical plan metrics. Since the majority of the bars show positive values, we can further confirm that GPT-RadPlan achieves greater OAR sparing in both prostate and head & neck plans.
Figure 6: Relative difference between GPT-RadPlan and clinical prostate plans. For every relevant OAR, each plot depicts the relative difference between clinical and GPT-RadPlan plans for 3 different metrics (mean dose, $D_{50}$ and $V_{15}$. Each white and gray block (each pair of bars) shows the results for a separate plan. Positive values indicate GPT-RadPlan superiority over clinical plans.
Figure 7: Relative difference between GPT-RadPlan and clinical heand and neck plans. For every relevant OAR, each plot depicts the relative difference between clinical and GPT-RadPlan plans for 3 different metrics (mean dose, $D_{50}$ and $V_{15}$. Each white and gray block with 6 bars shows the results for each of the plans individually. Positive values indicate GPT-RadPlan superiority over clinical plans.
B.2 Autoplan comparison
To further demonstrate GPT-RadPlan’s superiority to Autoplan, we have included the same plots and comparison used for comparison with clinical plans. Figure 8 compares the average DVH curves for both planning methodologies, which indicate a slightly more homogeneous PTV coverage in GPT-RadPlan plans, and overall improved OAR sparing, except for the femoral heads in prostate plans. For head & neck plans, Autoplan results in lower volumes of the brainstem and spinal cord receiving doses in the 10-20Gy range, although the maximum doses are ultimately higher.
Figure 9 further compares individual relative differences between plans, demonstrating that GPT-RadPlan outperforms Autoplan in most situations as shown by the positive values above the red line, except for the brainstem and spinal cord as shown in Figure 9(b).
Figure 8: Average DVH comparison between GPT-RadPlan and Autoplan. Visual comparison of the average DVH lines across all prostate (left) and lung (right) patients. Solid lines represent GPT-RadPlan plans, while dashed lines indicate Autoplan plans.
(a) Prostate patients
(b) Head & neck patients
Figure 9: Differences between GPT-RadPlan and Autoplan plans. For every relevant OAR, each box displays the relative difference between plans obtained using the Autoplan and GPT-RadPlan algorithms. Each panel shows the relative difference between different OAR metrics, including (left) the mean dose, (center) the $D_{50}$ and (right) the $V_{15}$. Positive values above the red line indicate that GPT-RadPlan plans result in lower metric values, which is preferred.